OpenStudy (anonymous):

Write in simplest form: 2.5b - 3(-2b). A. -3.5b B. 2.5b + 6b C. 8.5b

OpenStudy (anonymous):

c

OpenStudy (anonymous):

2.5b+6b=b(2.5+6)=8.5b